∫ 4u^{\frac{3}{4}}du


1
Expert's answer
2020-09-13T18:20:03-0400

Solution. Find the integral


"\intop 4u^{\frac{3}{4}}du=\frac{16u^{\frac{7}{4}}}{7}+C"

where C is constant.

Answer.

"\intop 4u^{\frac{3}{4}}du=\frac{16u^{\frac{7}{4}}}{7}+C"